# Yogesh Kulkarni, Pooyan Fazli

### v1

@@ Introduction
We then apply a parabolic function to generate a U-shaped weight $w_{t}$:
@@
where $\lambda_{\text{TAS}}$ controls the curve’s amplitude. This function assigns the minimum weight of 1.0 to tokens at the midpoint of the sequence ($\tilde{t}=0.5$) and the maximum weight of $1.0+\lambda_{\text{TAS}}$ to tokens at the start and end ($\tilde{t}=0$ or $1$). These weights produce a token-specific shaped advantage: $A^{\text{TAS}}_{i,t}=w_{i,t}\cdot A_{i}$ where $i\in\{1,\dots,K\}$ indexes each of the $K$ candidate responses in the group. AVATAR’s loss is formed by replacing the standard GRPO advantage $A_{i}$ in Equation 2 with this shaped advantage.
A key challenge is that reasoning sequence lengths, $L_{i}$, vary across completions. Our TAS mechanism is robust to this variability because the normalization step in Equation 5 makes the weight for any token $t$ depend on its relative position $\tilde{t}$ within the sequence, rather than its absolute position. This length-agnostic approach is supported by our ablation studies (Table 4), which show that our U-shaped weighting consistently outperforms the uniform baseline.
Reward Functions
@@
(1) Format Reward ($R_{\text{format}}$).
@@
$R_{\text{format}}(o_{i})=\begin{cases}1,&\text{if format is correct}\\ -1,&\text{otherwise}\end{cases}$ (7)
(2) Final Answer Accuracy ($R_{\text{acc}}$).
A black-box evaluation of the final answer $a_{i}$ extracted from the <answer> tag, compared to the ground truth $a_{\text{GT}}$. For numerical tasks, we use the relative Mean Absolute Error (rMAE) for a dense reward signal (Lu et al. 2025).
$R_{\text{acc}}(a_{i})=1.0-\min\left(1.0,\frac{|a_{i}-a_{\text{GT}}|}{a_{\text{GT}}}\right)$ (8)
(3) Self-Rewarding ($R_{\text{self}}$).
We use self-rewarding to help the model learn from its own consensus. Given a group of $K$ generated answers $\{o_{1},o_{2},\dots,o_{K}\}$ for a prompt, we choose the most common answer as a “pseudo-correct” $o^{*}_{\text{maj}}$ one using majority vote.
$o^{*}_{\text{maj}}=\text{argmax}_{o\in\{o_{1},\dots,o_{K}\}}\sum_{i=1}^{K}\mathbb{I}(o_{i}=o),$ (9)
@@
$R_{\text{self}}(o_{i})=\begin{cases}1,&\text{if }o_{i}=o^{*}_{\text{maj}}\\ 0,&\text{otherwise}\end{cases}$ (10)
(4) Stepwise Reasoning Judge ($R_{\text{judge}}$).
@@
A white-box reward from a frozen VLM judge (InternVL-3B (Zhu et al. 2025)), which provides a detailed score for reasoning quality within the <think> block. It evaluates logical consistency and correct use of grounding clues from the prompt. The score $s_{\text{judge}}\in[0,1]$ reflects the reasoning process quality independently of the final answer.
Video-Context Reference Score (VCRS).
To ensure a stable learning signal, we introduce the Video-Context Reference Score (VCRS), named for its role as a stable reference score based on the historical context of a specific video prompt. VCRS acts as a multiplicative factor in the final advantage calculation. It is computed as a moving average of rewards over the last 20 processed instances of a given prompt, ensuring the advantage baseline is never zero and helping to prevent learning stagnation on difficult tasks.
We apply AVATAR within a three-stage RL training pipeline (Figure 3) to evaluate its effectiveness. Stage 1 develops general visual reasoning, Stage 2 introduces audio-visual alignment, and Stage 3 addresses fine-grained audio-based object localization. Each stage leverages specific reward functions to ensure a stable and informative learning signal: Stage 1 uses format and accuracy rewards, Stage 2 augments these with the self-rewarding mechanism, and Stage 3 combines format rewards with the stepwise reasoning judge.
Audio-Visual Benchmarks
Model Reasoning OmniBench DailyOmni AV-Counting AV-Odyssey WorldSense IntentBench
Qwen2.5 Omni $\times$ 44.2 44.0 22.3 29.8 44.2 63.7
Echolnk $\checkmark$ 46.5 46.2 22.7 31.1 45.7 63.6
Omni-R1 $\checkmark$ 46.9 46.8 22.0 31.2 44.1 63.5
HumanOmni $\checkmark$ 44.9 47.6 19.6 30.3 45.4 *
Ola-7B $\times$ 45.3 52.3 17.4 25.6 44.2 59.1
AV-Reasoner $\checkmark$ 48.3 53.8 23.0 25.6 44.6 59.5
AVATAR $\checkmark$ 49.1+4.9 47.0+3.0 23.1+0.8 32.1+2.3 46.0+1.8 63.9+0.2
Table 1: Comprehensive evaluation of AVATAR vs. state-of-the-art audio-video understanding models. Best scores are in bold, and second-best scores are underlined. For AVATAR, performance improvements over Qwen-2.5-Omni are shown as subscripts. All results are reproduced by us and * refers to data contaminated.
General Video Understanding Video Reasoning
Model Reasoning MVBench Video-MME LVBench Video-Holmes MMVU TOMATO
Omnimodal Baselines (with Audio)
Qwen2.5 Omni $\times$ 66.1 58.3 37.2 40.6 60.2 29.0
Echolnk $\checkmark$ 66.2 60.8 37.6 42.5 65.7 29.9
Omni-R1 $\checkmark$ 66.0 60.7 37.6 44.2 63.6 29.2
HumanOmni $\checkmark$ 61.4 63.1 36.2 39.6 61.8 27.1
Ola - 7B $\times$ 40.1 59.1 35.5 40.1 56.6 -
AV-Reasoner $\checkmark$ 47.9 56.8 33.7 39.6 57.9 -
Video-only Reasoning Baselines (on Qwen2.5-VL)
Qwen2.5-VL $\times$ 59.4 60.1 31.2 38.0 61.9 29.3
Video-R1 $\checkmark$ 63.9 59.3 27.8 41.0 63.1 19.8
VideoRFT $\checkmark$ 62.1 59.8 34.2 40.5 58.0 20.6
TW-GRPO $\checkmark$ 63.3 60.1 33.8 38.4 61.4 23.5
AVATAR $\checkmark$ 66.4+0.3 62.8+4.5 38.4+1.2 45.1+4.5 65.6+5.4 30.8+1.8
Table 2: Comprehensive evaluation of AVATAR on general video understanding and reasoning benchmarks. Best scores are in bold, and second-best scores are underlined. For AVATAR, performance improvements over Qwen-2.5-Omni are shown as subscripts. All results are reproduced by us.
Model Configuration Audio-Visual Video Reasoning
OmniBench DailyOmni AV-Odyssey Video-Holmes MMVU TOMATO
(1) Baseline Qwen2.5 Omni 44.2 44.0 29.8 40.6 60.2 29.0
(2) Baseline GRPO 45.4+1.2 44.8+0.8 31.3+1.5 43.2+2.6 64.0+3.8 29.2+0.2
+ TAS Only 45.1+0.9 45.4+1.4 31.4+1.6 43.4+2.8 64.3+4.1 29.5+0.5
+ Curriculum & Replay 47.8+3.6 45.9+1.9 31.6+1.8 43.7+3.1 64.8+4.6 29.3+0.3
AVATAR (w/o Hinting) 48.2+4.0 46.1+2.1 31.8+2.0 44.3+3.7 65.0+4.8 30.2+1.2
AVATAR (w/ Hinting) 49.1+4.9 47.0+3.0 32.1+2.3 45.1+4.5 65.6+5.4 30.8+1.8
Table 3: Component-wise ablation study across audio-visual and video reasoning benchmarks.
Experiments

### v2

@@ 4.2 Temporal Advantage Shaping (TAS)
We then compute a token-specific weight using a parabolic function,
@@
where $\lambda_{\text{TAS}}$ controls the amplitude of the weight variation. This function assigns the minimum weight of 1.0 to tokens at the midpoint of the sequence ($\tilde{t}=0.5$) and the maximum weight of $1.0+\lambda_{\text{TAS}}$ to tokens at the start and end ($\tilde{t}=0$ or 11). These weights define a token-specific shaped advantage: $A^{\text{TAS}}_{i,t}=w_{i,t}\cdot A_{i}$, where $i\in\{1,\dots,K\}$ indexes each of the $K$ candidate responses in the group. AVATAR’s loss is formed by replacing the standard GRPO advantage $A_{i}$ in Equation 2 with this shaped advantage. Reasoning sequence lengths, $L_{i}$, vary across completions, which can affect positional weighting. However, the TAS mechanism remains robust because the normalization in Equation 5 ensures that each token’s weight depends on its relative position $\tilde{t}$ within the sequence rather than its absolute position.
4.3 Reward Functions
@@
(1) Format Reward ($R_{\mathrm{format}}$).
@@
$R_{\mathrm{format}}(o_{i})=\begin{cases}1,&\text{if format is correct}\\ -1,&\text{otherwise}\end{cases}$ (7)
(2) Final Answer Accuracy ($R_{\mathrm{acc}}$).
A black-box evaluation of the final answer $a_{i}$ extracted from the <answer> tag, compared to the ground truth $a_{\text{GT}}$. For non-zero numerical tasks, we use the relative Mean Absolute Error (rMAE) for a dense reward signal [avreasoner].
$R_{\mathrm{acc}}(a_{i})=1.0-\min\left(1.0,\frac{|a_{i}-a_{\text{GT}}|}{a_{\text{GT}}}\right)$ (8)
(3) Self-Rewarding ($R_{\mathrm{self}}$).
We use self-rewarding to help the model learn from its own consensus. Given a group of $K$ generated answers $\{o_{1},o_{2},\dots,o_{K}\}$ for a prompt, we choose the most common answer as a “pseudo-correct” $o^{*}_{\text{maj}}$ one using majority vote. Ties are broken by choosing the response with the highest model confidence (average token likelihood).
$o^{*}_{\mathrm{maj}}=\mathrm{argmax}_{o\in\{o_{1},\dots,o_{K}\}}\sum_{i=1}^{K}\mathbb{I}(o_{i}=o),$ (9)
@@
$R_{\mathrm{self}}(o_{i})=\begin{cases}1,&\text{if }o_{i}=o^{*}_{\mathrm{maj}}\\ 0,&\text{otherwise}\end{cases}$ (10)
(4) Stepwise Reasoning Judge ($R_{\mathrm{judge}}$).
@@
A white-box reward from a frozen VLM judge (InternVL3 [internvl3]), which provides a detailed score for reasoning quality within the <think> block. It evaluates logical consistency and correct use of grounding clues from the prompt. The score $s_{\text{judge}}\in[0,1]$ reflects the reasoning process quality independently of the final answer. The judge prompt is provided in Supplementary, Figure 14.
4.4 Video-Context Reference Score (VCRS)
Standard off-policy updates can be noisy, as the group mean ($\mu_{R}$) is an unstable baseline for stale buffer samples. We stabilize learning by introducing VCRS, denoted as $\bar{R}(q)$, which is a stable moving average of rewards for a prompt $q$ computed over its last 20 new (on-policy) instances.
This stable score $\bar{R}(q)$ replaces the noisy $\mu_{R}$ in the off-policy advantage calculation:
$A_{i,\text{off}}=\frac{R(o_{i})-\bar{R}(q)}{\sigma_{R,\text{off}}},$ (11)
where $\sigma_{R,\text{off}}$ is the standard deviation of the off-policy batch. This provides a reliable advantage signal, prevents stagnation, and makes the GRPO process more robust. This final advantage $A_{i,\text{off}}$ is then modulated by TAS.
Table 1: Dataset and Reward Configuration for AVATAR Training.
Stage Focus Datasets Reward Components
S0: SFT Cold Start Video-R1-CoT [videor1], TPO [tpo] N/A (Supervised)
S1: RL Visual Reasoning Video-R1 [videor1] $0.5\times R_{\text{format}}+0.5\times R_{\text{acc}}$
S2: RL Audio-Visual Reasoning AVQA [avqa], AVE [ave] $0.2\times R_{\text{format}}+0.4\times R_{\text{acc}}+0.4\times R_{\text{self}}$
S3: RL Audio-Object Localization AVSBench [avsbench] $0.2\times R_{\text{format}}+0.4\times R_{\text{acc}}+0.4\times R_{\text{judge}}$
Table 2: Comprehensive evaluation of AVATAR vs. state-of-the-art audio-video understanding models. Best scores are in bold, and second-best scores are underlined. Performance improvements from applying AVATAR to a baseline model are shown in green with 95% confidence interval (CI) margins ($\pm$) computed via bootstrap resampling (10,000 iterations, percentile method). Improvements are statistically significant at $\alpha=0.05$ when CIs exclude zero. All results were reproduced by us; * denotes potential data contamination.
Model OmniBench DailyOmni AV-Counting AV-Odyssey WorldSense IntentBench State-of-the-Art Models Echolnk [echoink] 46.5 46.2 22.7 31.1 45.7 63.6 Omni-R1 [omnir1] 46.9 46.8 22.0 31.2 44.1 63.5 HumanOmni [humanomni] 44.9 47.6 19.6 30.3 45.4 * AV-Reasoner [avreasoner] 48.3 53.8 23.0 25.6 44.6 59.5 Model-Agnostic Reinforcement Learning with AVATAR Ola-7B [ola] (Baseline) 45.3 52.3 17.4 25.6 44.2 59.1     + AVATAR[Uncaptioned image] 47.2 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+1.9}\pm\textbf{0.5})}$ 55.7 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+3.4}\pm\textbf{0.7})}$ 19.5 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+2.1}\pm\textbf{0.6})}$ 28.8 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+3.2}\pm\textbf{0.6})}$ 45.0 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+0.8}\pm\textbf{0.4})}$ 61.9 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+2.8}\pm\textbf{0.6})}$ Qwen2.5-Omni [qwen25omni] (Baseline) 44.2 44.0 22.3 29.8 44.2 63.7     + AVATAR[Uncaptioned image] 49.1 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+4.9}\pm\textbf{0.7})}$ 47.0 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+3.0}\pm\textbf{0.6})}$ 23.1 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+0.8}\pm\textbf{0.4})}$ 32.1 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+2.3}\pm\textbf{0.6})}$ 46.0 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+1.8}\pm\textbf{0.5})}$ 63.9 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+0.2}\pm\textbf{0.1})}$
Table 3: Comprehensive evaluation of AVATAR on video understanding and reasoning benchmarks. Best scores are in bold, and second-best scores are underlined. Performance improvements from applying AVATAR to a baseline model are shown in green with 95% confidence interval (CI) margins ($\pm$) computed via bootstrap resampling (10,000 iterations, percentile method). Improvements are statistically significant at $\alpha=0.05$ when CIs exclude zero. Improvement not statistically significant ($p>0.05$; 95% CI includes zero).
General Video Understanding Video Reasoning Model MVBench Video-MME LVBench Video-Holmes MMVU TOMATO State-of-the-Art Models Echolnk [echoink] 66.2 60.8 37.6 42.5 65.7 29.9 Omni-R1 [omnir1] 66.0 60.7 37.6 44.2 63.6 29.2 HumanOmni [humanomni] 61.4 63.1 36.2 39.6 61.8 27.1 AV-Reasoner [avreasoner] 47.9 56.8 33.7 39.6 57.9 24.9 Qwen2.5-VL [qwen25vl] 59.4 60.1 31.2 38.0 61.9 29.3 Video-R1 [videor1] 63.9 59.3 27.8 41.0 63.1 19.8 VideoRFT [videorft] 62.1 59.8 34.2 40.5 58.0 20.6 TW-GRPO [twgrpo] 63.3 60.1 33.8 38.4 61.4 23.5 Model-Agnostic Reinforcement Learning with AVATAR Ola-7B [ola] (Baseline) 40.1 59.1 35.5 40.1 56.6 25.3      + AVATAR[Uncaptioned image] 45.4 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+5.3}\pm\textbf{0.8})}$ 61.4 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+2.3}\pm\textbf{0.6})}$ 36.6 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+1.1}\pm\textbf{0.5})}$ 42.4 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+2.3}\pm\textbf{0.6})}$ 57.3 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+0.7}\pm\textbf{0.4})}$ 26.6 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+1.3}\pm\textbf{0.5})}$ Qwen2.5-Omni [qwen25omni] (Baseline) 66.1 58.3 37.2 40.6 60.2 29.0      + AVATAR[Uncaptioned image] 66.4 $\scriptstyle{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}(\textbf{+0.3}\pm\textbf{0.3})}^{\ddagger}$ 62.8 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+4.5}\pm\textbf{0.7})}$ 38.4 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+1.2}\pm\textbf{0.5})}$ 45.1 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+4.5}\pm\textbf{0.7})}$ 65.6 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+5.4}\pm\textbf{0.8})}$ 30.8 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+1.8}\pm\textbf{0.5})}$
Table 4: Component-wise ablation demonstrating how each component addresses specific GRPO limitations. The stratified replay buffer resolves data inefficiency and vanishing advantages, TAS improves credit assignment, and hinting helps escape local optima.
Audio-Visual Benchmarks Video Reasoning Benchmarks Model Configuration OmniBench DailyOmni AV-Odyssey Video-Holmes MMVU TOMATO Qwen2.5 Omni 44.2 44.0 29.8 40.6 60.2 29.0 GRPO 45.4 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+1.2})}$ 44.8 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+0.8})}$ 31.3 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+1.5})}$ 43.2 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+2.6})}$ 64.0 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+3.8})}$ 29.2 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+0.2})}$ + TAS Only (w/ On-Policy GRPO) 45.1 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+0.9})}$ 45.4 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+1.4})}$ 31.4 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+1.6})}$ 43.4 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+2.8})}$ 64.3 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+4.1})}$ 29.5 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+0.5})}$ + Replay Buffer (w/ Uniform Credit) 47.8 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+3.6})}$ 45.9 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+1.9})}$ 31.6 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+1.8})}$ 43.7 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+3.1})}$ 64.8 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+4.6})}$ 29.3 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+0.3})}$ AVATAR (w/o Hinting) 48.2 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+4.0})}$ 46.1 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+2.1})}$ 31.8 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+2.0})}$ 44.3 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+3.7})}$ 65.0 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+4.8})}$ 30.2 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+1.2})}$ AVATAR (w/ Hinting) 49.1 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+4.9})}$ 47.0 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+3.0})}$ 32.1 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+2.3})}$ 45.1 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+4.5})}$ 65.6 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+5.4})}$ 30.8 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+1.8})}$
5 Experiments